### v1

# nonblind {nonblind}

Abstract
Algorithms frequently assist, rather than replace, human decision-makers. However, the design and analysis of algorithms often focus on predicting outcomes and do not explicitly model their effect on human decisions. This discrepancy between the design and role of algorithmic assistants becomes of particular concern in light of empirical evidence that suggests that algorithmic assistants again and again fail to improve human decisions. In this article, we formalize the design of recommendation algorithms that assist human decision-makers without making restrictive ex-ante assumptions about how recommendations affect decisions. We formulate an algorithmic-design problem that leverages the potential-outcomes framework from causal inference to model the effect of recommendations on a human decision-maker’s binary treatment choice. Within this model, we introduce a monotonicity assumption that leads to an intuitive classification of human responses to the algorithm. Under this monotonicity assumption, we can express the human’s response to algorithmic recommendations in terms of their compliance with the algorithm and the decision they would take if the algorithm sends no recommendation. We showcase the utility of our framework using an online experiment that simulates a hiring task. We argue that our approach explains the relative performance of different recommendation algorithms in the experiment, and can help design solutions that realize human–AI complementarity.
00footnotetext: Bryce McLaughlin (brycem@stanford.edu) and Jann Spiess (jspiess@stanford.edu), Graduate School of Business, Stanford University. We thank Sanmi Koyejo, Jason Weitze, Emma Rockall, Adrienne Propp, Carl Meyer, Omer Shiran-Cohen, and audience members at the 2023 INFORMS Annual Meeting and at Stanford for helpful comments and suggestions.
1 Introduction
Algorithms often assist, rather than replace, human decision-makers. For example, in the United States, federal pretrial service officers set bail for defendants after observing a Pretrial Risk Assessment (PTRA), rather than algorithmically implementing a bail decision in response to the PTRA. Similarly, doctors may be assisted by clinical decision support systems throughout the patient care process (Sutton et al., 2020) and managers may be assisted by pre-employment assessments throughout the hiring process (Raghavan et al., 2020). In all of these cases, the human maintains the final decision authority. However, the design and analysis of these algorithms often focus on the prediction of outcomes as if algorithmic recommendations were implemented directly, rather than how these recommendations affect human decisions.
The narrow focus of existing approaches on predicting outcomes rather than optimizing for human–AI complementarity becomes of particular concern in light of a growing body of evidence that shows that human decision-makers with access to algorithmic recommendations often make worse decisions than if these recommendations had been implemented directly, and in some cases may even take worse decisions than they would have taken without access to the algorithm. In a recent review, Hemmer et al. (2021) identifies 53 articles that report (i) unassisted human performance, (ii) algorithm performance, and (iii) algorithm-assisted human performance on a decision task. Only in 16 of these does the algorithm-assisted human decision-maker perform best. In many of these studies, decision-makers exhibit noisy and biased adoption of the algorithm’s recommendations, which counteracts the gains associated with the additional information the algorithm provides. This suggests the necessity to incorporate decision-makers’ adoption decisions explicitly when designing the algorithm.
This article formalizes the design of recommendation algorithms that assist humans in making binary treatment choices via an optimization problem that leverages the potential outcomes framework from causal inference. For the analysis of recommendation algorithms, we employ the notion of “potential outcomes” and make connections to the analysis of instrumental variables in causal inference with imperfect compliance. This view suggests that an optimal recommendation algorithm should take into account the final outcome of interest to which the decision is related as well as the decision-maker’s choice of when to comply with a recommendation. We then show that this framework is helpful in designing better recommendation algorithms and making sense of their empirical performance using an online experiment.
In Section 2, we introduce a principal–agent model that formalizes the design of recommendation algorithms that assist a human in making binary treatment choices. As a motivating example, we consider a manager making hiring decisions on applicants who are evaluated by a pre-employment assessment. The principal in our problem is the company, who designs an algorithm that assists a hiring manager (the agent) in making the ultimate hiring decision. This model allows for complementarity in information and expertise between the principal’s algorithm and the agent, thus moving beyond models in which perfect adherence to the algorithmic recommendation is optimal.
Section 3 connects our problem setup to the potential-outcomes framework from causal inference. There, we introduce a monotonicity assumption that allows us to describe the effect of recommendations in terms of compliance, similar to the analysis of instrumental variables by Imbens and Angrist (1994). The potential-outcomes framework formalizes how individual outcomes or actions change in response to treatment choices. We use this framework to capture how different recommendations given by the algorithm lead to different decisions by the human decision-maker. Using our potential outcomes, we can state the optimal recommendation algorithm as minimizing an objective that separates into two parts, (1) the loss incurred by a direct implementation of the algorithm’s recommendations and (2) the loss incurred (or avoided) due to how decision-makers deviate from the recommendations. Our monotonicity assumption then implies that recommendations can only move the decision-maker toward the recommended action.
In Section 4, we showcase our framework by applying it to an online experiment that involves a hypothetical hiring task. This experiment asks subjects to make 25 hiring decisions on hypothetical applicants, and randomly varies the recommendation algorithm across subjects. This setup allows us to evaluate how the structure of recommendations impacts ultimate task performance. Specifically, the separate recommendation algorithms are based on different assumptions about compliance behaviors. In line with our framework, we find that subjects perform better when the recommendations are designed to provide complementary information, rather than optimized for hiring decisions directly. The best-performing algorithm assumes that human decision-makers comply with the algorithm only selectively in those cases where they are uncertain. The resulting decisions outperform unassisted human decisions as well as the best performance by an algorithm on its own. We thus demonstrate that our framework can help design systems that improve decisions through human–AI complementarity.
Our work responds to an empirical literature that analyzes inefficiencies in the interactions between human decision-makers and prediction algorithms. This literature has identified cases where human decision-makers imperfectly use prediction algorithms in high-stakes decisions in criminal justice (Imai et al., 2021; Stevenson and Doleac, 2022; Angelova, Dobbie, and Yang, 2023), healthcare (Agarwal et al., 2023; Maron et al., 2020), and child protective services (Fogliato, De-Arteaga, and Chouldechova, 2022; Mills and Grimon, 2022), so much so that the assisted human decisions are often no better than unassisted human decisions. A more extensive review of this literature is detailed by Lai et al. (2021). Of the works in this literature, our empirical approach is most related to that of Green and Chen (2019) who posit ways human decision-makers should respond to algorithms and test these hypotheses in the lab. Relative to this literature, we focus on designing complementary recommendations, rather than focusing on describing ways in which humans misuse algorithms.
We join a collection of works spanning the areas of operations, economics, and computer science that design algorithms to assist humans in taking better decisions. Bansal et al. (2019) and Donahue, Chouldechova, and Kenthapadi (2022) propose improvements to collaborative decision processes based on the private information of each agent. Balakrishnan, Ferreira, and Tong (2022) and Orfanoudaki et al. (2022) find that humans often incorrectly assume linear relationships, then develop strategies to help human–algorithm decision systems overcome this issue. Athey, Bryan, and Gans (2020) and McLaughlin and Spiess (2022) model how effort costs and reference points, respectively, affect a human’s adoption of algorithmic assistance. Baek et al. (2023) and Bastani, Bastani, and Sinchaisri (2022) apply ideas from reinforcement learning to incorporate past human actions into the recommendations of assistive algorithms. Caro and de Tejada Cuenca (2023) and Sun et al. (2022) use data on workers’ deviations from algorithm advice to increase compliance. Noti and Chen (2023) focuses on the question of when to provide assistance, and shows that strategically restricting advice can preserve complementary between machine and human performance. Our approach differs from these prior works as it provides a causal framework to capture how a human’s response varies with the underlying algorithm. Similar to the work of Ibrahim, Kim, and Tong (2021) (where human predictors assist algorithmic decision-makers), our results highlight the importance of private information in designing impactful collaboration.
Our methods build upon a literature that optimizes the design of estimators that inform decision-making. We combine models of communication between heterogeneous agents (Kamenica and Gentzkow, 2011; Andrews and Shapiro, 2021; Spiess, 2022) with approaches from causal inference with imperfect compliance (Angrist, Imbens, and Rubin, 1996; Vytlacil, 2002; Mogstad and Torgovitsky, 2018) in order to design optimal recommendation algorithms in the prescriptive sense (Bertsimas and Kallus, 2020). This approach parallels work that transfers concepts from causal inference for use in decision-making (Manski, 2021, 2023), in particular the approach in Saghafian (2023) of incorporating unobserved confounding using sets of causal models and the approach in Ben-Michael et al. (2022) of learning robust decision policies from non-stochastic treatment assignments.
2 Problem Setup
We introduce a principal–agent model of joint human–algorithm decision-making{nonblind}, which builds upon McLaughlin and Spiess (2022). In this model, the principal designs an algorithm that gives recommendations to an agent. The agent is a human decision-maker who observes each recommendation and subsequently makes a binary treatment choice. As our main illustration, we apply this model to the problem a company (principal) faces when implementing an algorithmic pre-employment assessment to assist managers (agents) in the hiring process. The company wants its managers to hire good workers and not hire bad workers. The company impacts the managers’ decisions only through the recommendation algorithm they implement. We allow for the managers to have both unknown information and preferences, allowing for complex relationships between the algorithm the company implements and the decisions that managers take.
2.1 Hiring as a Binary Treatment Choice Problem
We consider the problem of taking a binary decision $D$ before a binary outcome $Y$ is realized. We capture the impact of this decision through a loss function $\ell$, which describes regret from implementing $D$ compared to the optimal decision. For each problem instance, we assume access to characteristics $X$ that have predictive power for $Y$.
As a running example throughout this paper, as the binary treatment choice we consider a hiring decision $D\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$ with the outcome $Y\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} G}}\}$ measuring the applicant’s job performance, which is unknown at the time of the decision. We measure the utility associated with this decision via the loss function,
$\ell(Y,D)=c_{I}\>\mathbf{1}[Y=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} G}},D=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}]+c_{II}\>\mathbf{1}[Y=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} B}},D=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}].$ (1)
The type-I error cost $c_{I}>0$ captures the value of a good ( G) worker’s productivity minus their hiring cost while the type-II error cost $c_{II}>0$ captures the hiring cost minus whatever productivity a bad ( B) worker provides.
While worker performance $Y$ is ex-ante unknown, we assume that there is information $X$ collected about the worker which is available at the time $D$ is made. We do not assume any specific structure of $X$, as the data available about applicants before hiring may include their resume, written question responses, video responses to interview questions, or the actions the worker took during a game (Li et al., 2021). We do require that the information $X$ is collected, documented, and made available prior to making a hiring decision $D$ on each worker, and for simplicity assume that $X$ has finite support.
2.2 Algorithm Design as a Principal–Agent Problem
We examine the decision $D$ through the lens of a principal–agent model with asymmetric information and potentially misaligned preferences. Specifically, we assume the decision $D$ is taken by an agent who (1) observes private information $U$, (2) does not observe characteristics $X$ directly, and (3) may not share the preferences represented by the loss function $\ell$. To improve the decisions by the agent, the principal (whose preferences are captured faithfully by $\ell$) implements an algorithm $f^{\textnormal{rec}}$ that provides recommendations $R=f^{\textnormal{rec}}(X)$ to the agent. Thus, similar to the information design literature (Kamenica and Gentzkow, 2011; Bergemann and Morris, 2019), the principal has partial control over the information the agent receives before they make the final decision.
Returning to our hiring example, we assume the principal is a company that implements an algorithmic recommendation system $f^{\textnormal{rec}}$ that uses the data $X$ on workers to generate hiring recommendations $R=f^{\textnormal{rec}}(X)$ which influence the hiring decisions of the agent, who is a manager at the company. The recommendations reduce the unstructured applicant data $X$ to a simpler form that is still informative about the worker’s ability $Y$. For simplicity, we limit the algorithm to either recommend a decision to the manager (not hire, $R=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}$, or hire, $R=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}$) or not make a recommendation at all ($R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$). In practice, recommendations may present more granular predictive information, such as a score that represents a scaled probability a worker is good ($Y=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} G}}$) and allows the managers to understand the algorithm’s relative ranking of different applicant workers. It may also include an explanation of the score, highlight pieces of $X$ the algorithm deems particularly relevant, or include certainty information about the confidence of its prediction (Raghavan et al., 2020).
The manager responds to the recommendation by making a hiring decision $D$ which may depend on the recommendation $R$ along with their private information $U$ and unknown private preferences. The manager may receive this private information $U$ as a result of interviews with the applicant and may have preferences not captured by $\ell$ which drive their hiring decisions. We do not, for now, assume any specific structure of the managers’ preferences or private information. Furthermore, similar to how we assume the company does not observe $U$, we assume the manager does not observe $X$. However, we allow arbitrary correlation between $X$ and $U$, effectively capturing the idea that $X$ and $U$ contain shared information. This flexibility allows us to capture the complex decisions managers exhibit empirically. Note that $D$ generally depends on the recommendation rule $f^{\textnormal{rec}}$ beyond the actual recommendation $R$ that is given in a specific case. This is because interpreting the recommendation $R$ requires an understanding of which cases lead to each recommendation.
Throughout, we include the option that the algorithm does not make a recommendation, $R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$, forcing the manager to make an active choice. We believe that this design feature has formal advantages for modeling recommendations and practical relevance for achieving human–algorithm complementarity. First, allowing for instances without recommendations enables us to include decisions taken without any recommendation algorithm (i.e., unassisted decisions) as a special case in our framework. Second, {nonblind} building upon McLaughlin and Spiess (2022), forcing active choices by avoiding recommendations in specific cases may reduce distortions in cases where human decision-makers know better. These beliefs are further supported by the work of Noti and Chen (2023), which finds that partially withholding algorithmic assistance (i) increases humans’ reliance on advice when given and (ii) can be leveraged strategically to improve humans’ prediction accuracy.
3 Making Sense of Recommendations with Potential Outcomes
Having set up a principal–agent problem for thinking about the design of recommendation algorithms, we now use the potential-outcomes framework from causal inference to discuss how the agent’s decision depends on the recommendation algorithm. Viewing the agent’s decision in this way suggests that the optimization of recommendations shares common features with the analysis of instrumental variables in inference problems. We strengthen this connection by applying a monotonicity assumption similar to that of Imbens and Angrist (1994), which allows us to represent the remaining potential outcomes in terms of the active decision the agent makes when no recommendation is given and whether the agent complies with the recommendation. This leads to a decomposition of the objective of the recommendation algorithm, in terms of compliance, into the performance of the recommended actions and the quality of active decisions.
3.1 Recommendations Induce Potential Outcomes
We now interpret the agent’s decision in terms of the potential-outcomes framework of Rubin (1974). Specifically, we write $D_{U}(r;f^{\textnormal{rec}})$ for the potential decisions that the agent would take when receiving the recommendation $r$ from the recommendation algorithm $f^{\textnormal{rec}}$. This notation makes explicit that potential outcomes vary with the private information $U$, and are specific to a recommendation rule $f^{\textnormal{rec}}$. The realized decisions are then obtained as $D=D_{U}(R;f^{\textnormal{rec}})$, where $R=f^{\textnormal{rec}}(X)$.
In our hiring example, the algorithm’s recommendations impact the company’s utility only through managers’ decisions. Our framework, therefore, parallels instrumental variables that affect outcomes only through their impact on a treatment variable. Rather than directly optimizing for the recommendation, the company must understand how the managers’ decision-response $D_{U}(r;f^{\textnormal{rec}})$ varies across $r\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$. Since decisions are themselves binary and the recommendations take three values, we can enumerate these potential outcomes and categorize them into four response types based on how the recommendation impacts the assisted managers’ decisions. These response types are similar in nature to the types developed by Angrist, Imbens, and Rubin (1996) for the analysis of instrumental variables.
Table 1 enumerates the potential decisions associated with a manager’s response to the recommendations and provides such a categorization. If the manager maintains the same decision regardless of the recommendation, they Ignore the recommendation. If they adopt the recommendation regardless of whether the algorithm recommends not hiring (N) or hiring ( H) the applicant, then they Comply with the recommendation. If, instead, the manager always takes the decision that the algorithm does not recommend, then they Defy the recommendation. If none of these cases hold, then it must be the case that the manager will Change their decision in response to the algorithm sending a recommendation, taking one decision when the algorithm sends no recommendation and taking the opposite decision if either recommendation is sent. While we maintain all these response types for now, we will later simplify the analysis by ruling out the Defy and Change types, analogously to the monotonicity assumption of Angrist, Imbens, and Rubin (1996).
Response type $D_{U}(\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}})$ $D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}})$ $D_{U}({\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}})$
Ignore N N N
H H H
Comply N N H
N H H
Defy H N N
H H N
Change H N H
N H N
Table 1: Enumeration and categorization of the potential outcomes associated with the managers’ decisions in response to different realizations of the recommendation. The Defy and Change response types will later be ruled out by 1.
3.2 Objective of Recommendation Algorithms
Having introduced a framework to capture potential decisions in response to recommendations, we now consider the optimal recommendation algorithm as the solution to an optimization problem that minimizes the loss achieved by the potential outcomes the algorithm induces. We show the optimal recommendation’s objective can be separated into a triage term that assumes that recommendations are implemented directly as well as an objective that captures the agent’s endogenous response to the recommendation. We compare this optimal recommendation algorithm to solutions that instead directly target optimal decisions, and provide an intuition for why that optimal decision rule will rarely be the optimal recommendation rule.
In our hiring example, the optimal recommendation algorithm for the company to implement is the mapping from applicant worker data to recommendations that minimizes the expected loss of the decisions taken by the manager,
$f^{\textnormal{rec}*}=\min_{f:X\mapsto R}\mathbb{E}\left[\ell(Y,D)\right]\text{ where }R=f(X),D=D_{U}(R;f).$ (2)
This optimization problem is quite difficult to solve, given the uncertainty about how managers will respond to recommendations. As a result, companies often implement algorithms that instead recommend the optimal decision based on the information available to the algorithm,
$f^{\textnormal{dec}*}=\min_{f:X\mapsto R}\mathbb{E}\left[\ell(Y,R)\right]\text{ where }R=f(X),$ (3)
implicitly assuming that the recommendation $R=f(X)$ gets implemented directly ($D=R$). An extension suggested in previous work by Raghu et al. (2019) is to triage instances into those where algorithms outperform humans and those where humans outperform algorithms. If we write $D_{0}=D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}})$ where $f^{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\equiv{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$ for the decision taken by the manager in absence of a recommendation algorithm, and we assume that the manager takes this decision when not given a recommendation ($f(X)={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$) and otherwise follows the algorithm, then this objective and its solution correspond in our framework to
$f^{\textnormal{tri}*}=\min_{f:X\mapsto R}\mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D_{0})\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]\text{ where }R=f(X).$ (4)
This objective again implicitly assumes that recommendations are followed perfectly when given and that manager decisions are not affected by the recommendation algorithm when no recommendation is given ($R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$). In practice, the decision for $R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$ may itself be different from the decision $D_{0}$ without a recommendation algorithm, and we may specifically hope it to improve since additional information and effort may be available for those cases (Raghu et al., 2019; Athey, Bryan, and Gans, 2020). Note also that the case of recommending an optimal decision is a special case where the algorithm always makes a recommendation ($R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$), with which the agent complies.
We now compare these different objectives and solutions, and show that the optimal recommendation ($f^{\textnormal{rec}*}$) can be thought of as an augmentation of the optimal triage ($f^{\textnormal{tri}*}$) that also incorporates the compliance behavior of the manager. Specifically, the loss function that defines the optimal recommendation can be decomposed into the loss function of the triage problem (which we call the triage effect, TE) and the response effect (RE) from managers deviating from the triage decision,
$f^{\textnormal{rec}*}=\min_{f:X\mapsto R}$ $\overbrace{\mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D_{0})\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]}^{\textnormal{{TE} -- Triage Effect}}$ (5)
$+\underbrace{\mathbb{E}\left[\left(\ell(Y,D)-\ell(Y,R)\right)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\left(\ell(Y,D)-\ell(Y,D_{0})\right)\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]}_{\textnormal{{RE} -- Response Effect}}$
where $R=f(X),D=D_{U}(R;f)$.
In addition to the triage effect from (4), this decomposition has a second part that represents deviations from recommendations and unassisted decisions. It can represent both improvements (in cases where the human decision-maker uses their private information to override a suboptimal recommendation or adjusts their active decision based on their knowledge of the recommendation) as well as mistakes (when the manager goes against a correct hiring recommendation or takes an active decision that is worse than in the case without any recommendations).
The two components also highlight the challenge of learning good recommendations from data. The triage effect (TE) is straightforward to estimate, provided the company has access to historical data through which they can learn the joint distribution of covariates $X$, ability $Y$, and unassisted manager decisions $D_{0}$. On the other hand, the response effect depends on the potential outcomes associated with the managers’ decisions, which complicates its evaluation for any $f^{\textnormal{rec}}$ under which decision data has not already been collected. Understanding $(D_{U}(\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}};f^{\textnormal{rec}}),D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}}),D_{U}(\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}};f^{\textnormal{rec}}))$ may require more than simple experimental evaluation based on local changes in recommendations. These potential decisions may vary arbitrarily with changes to $f^{\textnormal{rec}}$. Solving for the optimal recommendation then requires understanding this distribution across potentially implementable recommendation algorithms $\mathcal{F}\subseteq\{f;f:X\mapsto R\}$. This makes designing an optimal recommendation algorithm an intractable task without additional assumptions (even when we just want to robustly outperform the unassisted decisions $D_{0}$).
$\mathcal{F}$$f^{\textnormal{tri}*}$$f^{\textnormal{rec}*}$$f^{\textnormal{dec}*}$
Figure 1: Representation of potential algorithms $\mathcal{F}$ that could be implemented in terms of the triage effect TE and the response effect RE. The optimal triage $f^{\textnormal{tri}*}$ minimizes the triage effect over $\mathcal{F}$, while the optimal recommendation rule $f^{\textnormal{rec}*}$ minimizes the sum $TE+RE$. The optimal decision rule $f^{\textnormal{dec}*}$ need not lie on the Pareto frontier that minimizes these two losses, highlighted in red.
The relationship between algorithms that solve for the different objectives above is illustrated in Figure 1. Plotting the set of possible algorithms ($\mathcal{F}$) according to their triage effect (TE) and response effect (RE) clarifies, in particular, the relationship between optimal triage and optimal recommendation. The optimal triage ($f^{\textnormal{tri}*}$) minimizes the triage effect without considering the response effect related to each algorithm. The optimal recommendation ($f^{\textnormal{rec}*}$) potentially incurs additional loss through the triage effect in order to reduce the loss incurred due to the response effect. In practice, this means the optimal recommendation may be sub-optimal if applied directly, instead relying on the manager’s ability to strategically override decisions inconsistent with their own private information ($U$) and preferences. However, both $f^{\textnormal{rec}*}$ and $f^{\textnormal{tri}*}$ lie on the Pareto frontier which simultaneously minimizes the triage and response effects. This does not have to be true of the optimal algorithmic decision rule $f^{\textnormal{dec}*}$, which generally incurs a higher triage effect TE than the optimal triage solution $f^{\textnormal{tri}*}$. This is because the triage solution $f^{\textnormal{tri}*}$ comes from minimizing expected triage loss (4) subject to a less strict constraint than the decision solution $f^{\textnormal{dec}*}$. Meanwhile, the response effect can be better or worse when applying the $f^{\textnormal{dec}*}$ or $f^{\textnormal{tri}*}$ rules.
3.3 Monotonic Responses and Compliance-Aware Recommendations
Above, we have proposed a potential-outcomes framework for potential agent decisions in response to recommendations, and used this framework to discuss the design of optimal recommendation algorithms in general. We now introduce assumptions that limit the potential outcomes associated with the agent’s decision to simplify the analysis of their endogenous response. We then use these simplifications to shed further light on the structure of optimal recommendations in terms of agent compliance.
In the context of the four response types introduced in the hiring example and repeated in Table 1, manager behavior consistent with the Ignore and Comply responses is easier to explain than behavior consistent with the Defy and Change responses. If the manager believes their private information $U$ is much better than the algorithm’s for some problem instance, they are likely to Ignore the recommendation. Conversely, if they believe the algorithm’s information dominates any additional information in $U$, they will Comply. Using recommendations as a threshold requirement for considering an applicant, as done by hiring managers interviewed by Li et al. (2021), also generates potential outcomes consistent with only Ignore and Comply responses. On the other hand, Defy and Change behaviors are counter-intuitive and typically need bad recommendations, perverse correlation between $X$ and $U$, or adversarial managerial incentives to be justified rationally. Consequently, we now make an assumption to eliminate the potential outcomes associated with Defy and Change responses, similar to the monotonicity assumption of Angrist, Imbens, and Rubin (1996) in the context of instrumental variables and its extensions to multivariate discrete treatments (including Lee and Salanié, 2018).
Assumption 1 (Monotonic response to $f^{\textnormal{rec}}$).
[…]
At a high level this assumption restricts the direction of the managers’ response to follow $R$. Hire recommendations ($R=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}$) can turn not-hire decisions into hire decisions ($D=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}\to D=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}H}}$) but cannot turn hire decisions into not-hire decisions ($D=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\not\to D=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}$). Not-hire recommendations ($R=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}$) do the exact opposite. This restriction exactly rules out the bottom four rows in Table 1. It can be thought of as a restriction on plausible recommendation rules $\mathcal{F}$. Specifically, it rules out that recommendation rules are used that would make sense to Defy (such as a rule that suggests hiring bad candidates) or that would make sense to observe the Change behavior for (such as a rule that only gives a recommendation for bad candidates and never gives one for good candidates).
When the manager’s decision behavior is limited by 1, we can fully describe their behavior for some specific recommendation rule $f^{\textnormal{rec}}$ in terms of their active decisions $D_{\varnothing}=D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}})$ which they exhibit when not given a recommendation, but the algorithm $f^{\textnormal{rec}}$ is implemented, along with their compliance behavior $C$:
Proposition 1 (Simplified potential outcomes).
Under 1, the potential outcome triplet $(D_{U}(\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}};f^{\textnormal{rec}}),D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}}),D_{U}(\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}};f^{\textnormal{rec}}))$ can be fully expressed by the active decision $D_{\varnothing}=D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}})\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$, which expresses the agent’s decision when not given a recommendation, and the compliance type $C\in\{\textit{Ignore}{},\textit{Comply}{}\}$, which specifies whether the agent adopts $R$ when the recommendation is presented.
Table 2 visualizes 1 by restating the remaining potential-decision types in terms of $(D_{\varnothing},\textit{Comply})$ once 1 is applied. The only information determining the decision $D$ outside of the recommendation $R$ is the active decision $D_{\varnothing}$ and whether the response type $C$ of the manager is Ignore or Comply. The proposition follows directly from an inspection of Table 1. For completeness, a formal proof is given in Appendix A.
$C$ $D_{U}(\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}})$ $D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}})$ $D_{U}({\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}})$
Ignore $D_{\varnothing}$ $D_{\varnothing}$ $D_{\varnothing}$
Comply N $D_{\varnothing}$ H
Table 2: Reduction of potential outcomes under 1. This organization acts as a visual representation of 1, showing how the values that the potential outcomes take reduce to the active decision $D_{\varnothing}\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}{},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}{}\}$ absent a recommendation and the compliance $C\in\{\textit{Ignore}{},\textit{Comply}{}\}$.
As an immediate consequence of the simple representation of response types in 1, we can decompose the objective of the recommendation algorithm in terms of compliance:
$f^{\textnormal{rec}*}=\min_{f:X\mapsto R}$ $\mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[C=\textit{Comply}\text{ and }R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D_{\varnothing})\>\mathbf{1}\left[C=\textit{Ignore}\text{ or }R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]$ (6)
where $R=f(X)$.
Similarly, this explicit notation allows us to shed further light on the response effect from (5) in terms of compliance and deviations between active decisions $D_{\varnothing}$ and unassisted decisions $D_{0}$:
$\textit{RE}=\mathbb{E}\left[\left(\ell(Y,D_{\varnothing})-\ell(Y,R)\right)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},C=\textit{Ignore}\right]+\left(\ell(Y,D_{\varnothing})-\ell(Y,D_{0})\right)\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\right]$ (7)
The first term in this decomposition expresses deviations from recommendations when they are ignored, while the second term captures the change in behavior for those instances where no recommendation is given that comes from the presence of the recommendation algorithm.
Even with the simplifications introduced by 1, it is not generally possible to learn the optimal recommendation algorithm from past data $(Y,X,D_{0})$ on job performance, characteristics, and unassisted human decisions. This is because neither the compliance behavior $C$ nor the impact of recommendations on active decisions $D_{\varnothing}$ can be learned from such limited data. However, we can now formulate assumptions on these unobserved quantities in order to derive optimal learnable recommendation algorithms in specific cases, thus providing a formal analog and extension of the discussion in Section 3.2:
  • $\circ$
    If compliance is perfect ($C\equiv\textit{Comply}$) and decisions in instances without recommendations are not affected ($D_{\varnothing}\equiv D_{0}$), then the optimal recommendation algorithm is the optimal triage solution $f^{\textnormal{tri}*}$ (which can be learned from data on $(Y,X,D_{0})$).
  • $\circ$
    If compliance is perfect ($C\equiv\textit{Comply}$), and recommendations are limited to be binary ($R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$) or human decisions are dominated by recommended actions (such as when $\mathbb{E}[\ell(Y,D_{\varnothing})|X]\geq\min_{r\in\{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}},\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}\}}\mathbb{E}[\ell(Y,r)|X]$ $X$-a.s. and for all $f^{\textnormal{rec}}\in\mathcal{F}$), then the optimal recommendation algorithm is the optimal decision algorithm $f^{\textnormal{dec}*}$ (which can be learned from data on $(Y,X)$). This applies, specifically, in the case with perfect compliance and no private information of the human decision-maker.
At the same time, more nuanced assumptions or data on compliance behavior may be necessary to achieve complementarity in practical cases. This may include the cases below:
  • $\circ$
    If recommendations do not affect compliance ($C$) or active decisions for instances where no recommendations are given ($D_{\varnothing}=D_{0}$), then the optimal recommendation can generally be seen as the optimal triage decision among compliers only, that is, it solves
    $\min_{f:X\mapsto R}\mathbb{E}\left[\ell(Y,R)\>\mathbf{1}\left[R\neq{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]+\ell(Y,D_{0})\>\mathbf{1}\left[R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}\right]\middle|C=\textit{Comply}\right]\text{ where }R=f(X).$
    However, unlike the case of full compliance, we now have to learn about compliance behavior in addition to active decisions.
  • $\circ$
    Even if the agent takes optimal decisions given their available information, we would require information beyond $(Y,X,D_{0})$, since the optimal recommendation would also depend on the structure and information content of the private signal $U$.
In our experiment below, we combine information about the point distribution of information $(Y,X,U)$ with assumptions on active decisions $D_{\varnothing}$ and compliance $C$ in order to solve for sensible recommendation algorithms.
4 Complementary Recommendations in an Experiment
We now apply our framework to the design and analysis of recommendation algorithms in a controlled trial. In a pre-registered online experiment, study subjects participate in an incentivized hiring game. [^1] The pre-analysis plan for this experiment can be found at www.socialscienceregistry.org/trials/11857. This empirical application showcases our framework’s ability to map realistic assumptions about agent behavior to the design of practical recommendations that realize complementarity between humans and algorithms.
In the experiment, we collect hypothetical hiring decisions while varying the structure of the recommendation algorithm at the subject level. We apply different assumptions on how subjects will behave within our framework to generate these recommendation algorithms. We find that subjects are able to improve their decisions with algorithms whose recommendations are focused on providing suggestions that are complementary to the subjects’ private information. Moreover, subjects given these focused recommendations perform better than subjects who see naive recommendations based on optimal algorithmic predictions, and outperform decisions by the human or algorithm alone.
4.1 Experimental Design
We implement a hiring game as part of an online experiment to test subjects’ performance when using recommendation algorithms in a setting where both subjects and algorithms have private information. Specifically, we label 25 hypothetical applicants as good or bad based on the role they are applying for (private information only available to experiment subjects) and their personality type (private information only available to the algorithms), and ask subjects to only hire good applicants. We measure how the accuracy of subjects’ hiring decisions changes in response to different recommendation algorithms (treatments) that vary at the subject level.
For each subject, the experiment proceeds as follows. Subjects are brought to a Qualtrics survey from the recruiting site Prolific and informed they will make hiring decisions on 25 hypothetical applicants, each of which are good or bad types (referred to as the applicant’s ability). They are informed that each applicant’s ability $Y$ is determined by the relationship of the applicant’s personality type $X$ (denoted by letters A–E) to their intended role $U$ (Engineering, Sales, or Communications), as shown in Figure 2. They then answer three comprehension questions about how ability $Y$ is determined and, conditional on answering correctly, are randomized into a treatment. All but one control condition give subjects a recommendation algorithm to assist them with making hiring decisions. Unlike the study subjects, these algorithms can access the applicants’ personality types, but not their role. Subjects are then asked three more comprehension questions about the specific algorithm they will interact with, but continue in the experiment regardless of their answers. Subjects then make 25 hiring decisions with access to the distribution of applicants as represented in Figure 2, the recommendations of their assigned algorithm, and an explanation of that specific recommendation algorithm. An example of a hiring decision is given in Figure 3. When making these hiring decisions, subjects know the applicant’s role, but are unaware of the applicant’s personality type beyond any information conveyed by the recommendation.
Figure 2: Summary of how the information available to the subject (the applicant’s role) and the information of the algorithm (the applicant’s personality type) relate to the applicant’s ability (bad or good). This information is shown to all participants in the experiment and is easily accessible throughout their decision–making process.
Figure 3: Sample hiring decision from the Predictive treatment.
Subjects are incentivized to hire good applicants and to not hire bad applicants. Specifically, as compensation for participating in our experiment, subjects receive \$2.04 for their time (estimated to be 12 minutes) and the chance to earn a \$2.00 bonus as a performance incentive. They earn this bonus if their hiring decision on a randomly selected applicant is correct (hire if good, not hire if bad). Note that this induces equal type I and type II losses ($c_{I}=c_{II}$) from the perspective of our subjects, and we analogously evaluate their performance using the percentage of correct decisions taken. Subjects see all 25 applicants in a random order to make their hiring decisions, but don’t learn whether they took correct decisions.
The joint distribution of applicant role, personality type, and ability pictured in Figure 2 is designed to vary subjects’ willingness to Comply with recommendations across applicant roles while allowing the recommendation to provide useful information cases the applicant role is relatively uninformative. When subjects review an applicant for an Engineering role, they know the worker is good. Subjects are much more uncertain for Sales applicants (60% of which are good) and Communications applicants (40% of which are good). The algorithm’s private information (personality type) varies significantly across the cases that subjects are uncertain, thus containing useful information to help them make their decisions. Moreover, both the subject and algorithm’s private information are of equal value in the sense that a rational agent with access to either the role or the personality type would take the correct decision $76\%$ of the time.
4.2 Treatments
The recommendation algorithm a subject receives represents the treatment in our experiment and is randomized at the subject level. In addition to four treatment arms representing different recommendation algorithms, there is a control arm in which no recommendation is given. These recommendation algorithms vary by the way they map personality types $X$ to recommended actions $R=f^{\textnormal{rec}}(X)\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$, where the recommendations follow our earlier example and correspond to not hiring, no recommendation, and hiring the applicant, respectively. We now describe the four different algorithms, and link them to the implicit assumptions about compliance and active decisions that motivate their design. Additionally, Table 3 lists the recommendation algorithms that subjects could be assigned to, alongside the recommendations that these algorithms provide across personality types.
Algorithm A B C D E
[…]
Predictive N H H H H
Complementary N H H H N
[…]
Control.
Subjects in this condition did not observe a recommendation algorithm. We use this treatment to collect information on the distribution of unassisted decisions subjects take.
Predictive.
The algorithm in this condition sends a hire recommendation if at least half of the applicants associated with the given personality type are good and sends a not-hire recommendation otherwise. This Predictive algorithm is the optimal decision an algorithm can take by itself in order to minimize overall errors, so it corresponds to $f^{\textnormal{dec}*}$ in the framework of Section 3. It would correspond to the optimal recommendation if compliance was perfect and active decisions were only leading to worse hiring decisions.
Complementary.
These subjects observe an algorithm that sends a hire recommendation if at least half of the Sales and Communications applicants associated with the given personality type are good and sends a not-hire recommendation otherwise. This algorithm ignores Engineering applicants on the assumption that subjects will perform effectively by themselves when asked whether to hire these applicants, and instead sends the optimal decision for the subpopulation of Sales and Communications applicants. It corresponds to the optimal recommendation if subjects do not comply with the recommendation for Engineering candidates, but comply perfectly for the remaining applicants.
Triage.
These subjects observe a version of the Predictive algorithm that implements a ‘safety check’, causing the recommendation not to send recommendations for some personality types. Specifically, the Triage algorithm only sends the Predictive algorithm’s recommendation of hire if at least half of the Communications applicants associated with the given personality type are good and it only sent the Predictive algorithm’s recommendation of not hire if at least half of the Engineering and Sales applicants associated with the given personality type are bad. As a consequence, this algorithm does not send recommendations for the B and C personality types (which are equivalent). The Triage algorithm is the optimal triage algorithm ($f^{\textnormal{tri}*}$) for a rational agent with the subjects’ private information.
Complementary Triage.
These subjects observe an algorithm that combines the logic of the Triage and Complementary algorithms. These subjects receive the Complementary algorithm’s recommendation of hire if at least half of the Communications applicants associated with the given personality type are good and receive the Complementary algorithm’s recommendation of not hire if at least half of the Sales applicants associated with the given personality type are bad. The Complementary Triage algorithm is the optimal algorithmic triage over the subpopulation of Sales and Communications applicants.
We now summarize the logic behind these algorithms by examining their implicit assumptions about the subjects’ compliance choices and active decisions. Specifically, Table 4 presents each of the algorithms as a solution to a specific combination of assumptions. For compliance choices, we distinguish between perfect compliance (that is, following the recommendation and only making an active decision when no recommendation is given) and selective compliance. By selective compliance, we mean that subjects always take an active decision for non-Engineering candidates, and comply otherwise. For active decisions, we distinguish between completely random decisions and sophisticated decisions. Sophisticated decisions assume that the subject uses their private role information effectively when they take active decisions, that is, hire Engineering and Sales candidates, which is optimal absent recommendations and across each of our treatment arms. For example, if compliance is perfect and decisions are sophisticated, then we assume that subjects follow all recommendations and correctly decide in the absence of a recommendation when they do not receive one.
Using these combinations of behaviors, we can interpret the above treatment arms as optimal recommendation algorithms. First, the Predictive algorithm is optimal with perfect compliance and random active decisions. Second, the assumptions under which the Triage algorithm is optimal are that compliance is perfect and active decisions are sophisticated. In this case, the subject follows the recommendation when it is given, and otherwise relies on their private information to take a decision. Third, the Complementary algorithm is optimal when subjects comply for non-Engineering candidates only, and take random decisions when no recommendation is sent. Fourth, the Complementary Triage recommendations achieve optimal outcomes when subjects do not comply for (and always hire) Engineering candidates, and effectively use their private information to take a sophisticated decision when they do not receive a recommendation. In theory, this treatment arm would achieve the best performance on average, with only one mistake (not hiring a type-E communications candidate).
Compliance $C$
Active decision $D_{\varnothing}$ Perfect Selective
Random Predictive Complementary
Sophisticated Triage Complementary Triage
Table 4: Optimal algorithms under different combinations of compliance choices and active decisons.
Although final recommendations only depend on the personality type $X$, we note that constructing these algorithms requires an understanding of the joint distribution of $X$ with the outcomes $Y$ as well as the role $U$. Whether such a joint distribution can be learned in practice depends on the type and quality of data that is available at the time of construction of an algorithm. Here, we focus instead on an analysis of the relative performance of different recommendation algorithms in a setting where we (and the study subjects) know the joint distribution of $(Y,X,U)$.
4.3 Data and Main Results
Before analyzing the results of the experiment, we describe the data collection process, final sample sizes for each treatment, and basic performance statistics. Among 1675 surveys begun, about 35% failed to answer at least one of the initial comprehension questions and were screened out, leaving us with 1083 subjects that completed our study. Subjects were split approximately equally between all arms except Control, which received half as many observations as the other treatments. Specifically, each of the four treatment arms contained approximately 240 participants who, like the approximately 120 Control subjects who did not get recommendations, each took 25 hiring decisions for a total of around 6,000 decision-responses per treatment and 3,000 for Control. Overall, subjects performed well in the experiment and within the estimated time frame. The average fraction of correct decisions was 78.1% across all subjects. The average duration of the experiment was around twelve minutes. However, both the performance (standard deviation of 13 percentage points) and duration (standard deviation of almost eight minutes) show significant heterogeneity across subjects.
We compare treatments by average subject performance, with the main results reported in Table 5. Subjects in the Predictive treatment, who received recommendations from an algorithm that was trained to predict outcomes, got around 76% of hiring decisions correctly, significantly above the around 70% in the Control group that was not supplied with any recommendations. Subjects in the Triage, Complementary, and Complementary Triage treatments were all designed to provide recommendations that complement human decisions, and on average outperformed subjects in the Predictive treatment. Hence, the data confirms that the algorithm designed to take optimal decisions by itself (given by the Predictive treatment) is not the optimal recommendation algorithm in this case. Across the treatments, subjects assigned to the Compliance algorithm perform best, with over 81% optimal decisions (although this performance is statistically indistinguishable from the Complementary Triage subjects). Overall, the results suggest that subjects tend to take smart compliance choices and frequently (and correctly) override the algorithm for Engineering candidates. At the same time, decisions are inefficiently noisy.
Treatment arm Fraction optimal decisions (%) (SE)
Control 69.5 (0.7)
Predictive 76.2 (0.5)
Complementary 81.6 (0.7)
Triage 78.8 (0.8)
Complementary Triage 80.1 (1.1)
Feasible optimum 96.0
Algorithm by itself 76.0
Random choice 50.0
Table 5: Average subject performance (through fractions of optimal hiring decisions) across treatments. Bootstrap standard errors are given and clustered at the subject level. In addition to subject performance in the experiment, the feasible optimum considers the best theoretically achievable performance by a combination of a recommendation algorithm with three levels and a rational agent, while we also consider the optimal performance of an algorithm by itself as well as the trivial benchmark of a completely random hiring choice based on the flip of a fair coin.
We further compare subject performance to useful reference points in Figure 4 (justified further in Wu et al., 2023). Throughout, we focus on the fraction of cases subjects are taking the optimal decision on an applicant (hire if good, not hire if bad). Within each treatment, we compare subject performance to the performance when following the algorithm perfectly (corresponding to the Triage Effect of Section 3.2), and plot the impact of subject deviations (Response Effect). We also provide the best achievable performance by an algorithm acting alone (to benchmark which treatments were able to outperform algorithmic automation) as well as the performance achievable by a rational agent acting with the subject’s information (to act as a performance upper bound). The decomposition reveals that subject performance tends to improve by deviating from the algorithm (at least in cases other than the Predictive arm, in which case subject performance is equivalent to perfectly following the algorithm). Interestingly, machine-assisted human performance is best for the somewhat simpler Complementary algorithm than for the theoretically optimal Complementary Triage. Overall, we think that the patterns in Figure 4 are in line with good compliance choices, noisy active decisions, and a penalty for overly complex recommendation algorithms.
Figure 4: Average subject performance (through fractions of optimal hiring decisions) across treatments. $95\%$ bootstrap confidence intervals are clustered at the subject level. In addition, the triage performance (where compliance is perfect and not sending a recommendation uses the decision of Control) is given for each treatment by a gray dot, while the Response Effect from Section 3.2 is shown by the vertical red dashed lines. The optimal algorithm performance is shown by the gray horizontal line. The optimal performance achievable by a rational agent with the subject’s knowledge (role and algorithmic recommendation) is given for each treatment arm as a black dot.
Our experiment documents complementarity between humans and algorithms when recommendations are designed in the right way. Subjects in the Triage, Complementary, and Complementary Triage treatments all performed significantly better (at a 5% significance level) than humans by themselves (Control) or the algorithm by itself, thus achieving human-AI complementarity, while subjects in the Predictive treatment could only match the performance human or algorithm could (theoretically) achieve by themselves.
4.4 Additional Analysis and Robustness Checks
To understand performance differences across treatment arms further, we separate subjects’ propensity to hire applicants across roles and personality types in Figure 5, and relate this behavior to the recommendations subjects receive. From the graph, we can identify that subjects’ unassisted decisions mostly match the average quality of candidates (which would be in line with a probability-matching model where subjects match the probability they hire an applicant with the probability they are good). Their compliance choices are generally sophisticated (in that they correctly override the algorithm for Engineering candidates), although somewhat random. Note that our framework predicts that the Complementary algorithm is optimal if subjects make somewhat noisy active decisions but make sophisticated compliance choices. This is in line with our empirical findings, where the Complementary algorithm is the best among those algorithms we tested and likely the optimal recommendation algorithm we could have implemented.
Figure 5: Frequency of experiment subjects taking a hire decision ($y$-axis) across the different combinations of treatments (color) and profiles ($x$-axis). Different markers are used to identify whether the subject received a hire recommendation ($\Uparrow$), did not receive a recommendation (--), or received a not-hire recommendation ($\Downarrow$). The optimal decision given oracle information is given by the location of the points on the lines $y=1$ (hire is optimal) and $y=0$ (not hire is optimal). The dashed lines give the frequency of an applicant being hired by a subject without access to recommendations (Control arm).
As an additional robustness check for our results in this section, we exploit additional comprehension questions we asked about the specific recommendation algorithms. Specifically, 376 of the 961 study subjects who were given a recommendation did not answer one of the comprehension questions correctly that was asked about their specific algorithm. We reproduce all analysis on both the sub-population of subjects who answered all recommendation comprehension questions correctly and the sub-population of subjects who answered at least one question wrong in Appendix B. Our results strengthen on the sub-population who answered all comprehension questions correctly, while, as expected, results are very noisy and largely insignificant on the sub-population of subjects who answered at least one question wrong. These results support conclusions by Bansal et al. (2019) that a subject’s understanding of an algorithm is vital to achieving complementarity.
5 Conclusion
This article formalizes the analysis and optimal design of recommendation algorithms within a principal–agent model that employs the potential-outcomes framework from causal inference. We draw similarities between the impact that recommendation algorithms have on human decisions in algorithmic design and the effect that instrumental variables have on individual treatment choices in inference problems. We use this connection to decompose the objective function of a recommendation algorithm into algorithmic performance and human response. Making additional assumptions on reasonable human responses to recommendations, our approach allows us to express recommendation-assisted decisions in terms of compliance and active decisions. We can then reason about optimal recommendation algorithms in terms of assumptions on each component individually. We demonstrate the utility of our framework in a controlled online experiment, which reinforces that decision-maker responses should be considered when designing recommendation algorithms. Specifically, we document that subjects perform better using recommendations that are designed to provide complementary information, relative to algorithms that are instead designed to solve a problem directly.
Appendix
Appendix A Proofs
Proof of 1.
To show this result we will build a bijection between the potential outcomes $(D_{U}(\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}};f^{\textnormal{rec}}),D_{U}({{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}};f^{\textnormal{rec}}),D_{U}(\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}};f^{\textnormal{rec}}))$ which exist under 1 and realizations of $(D_{\varnothing},C)$:
[…]
Appendix B Robustness Checks
Here we reproduce the results of Figure 4 and Figure 5 for the sub-populations who answered all of their recommendation comprehension questions correctly (585 subjects) and those who answered at least one question wrong (376 subjects). Subjects who answered at least one recommendation comprehension question incorrectly exhibited noisier decisions on average than those who answered all questions correctly. We find that algorithms generated using our framework exhibited higher levels of complementarity for the sub-population who correctly answered all comprehension questions while the sub-population who answered at least one comprehension question wrong failed to realize complementarity.
Figure 6 reports each sub-populations average performance by treatment. Both sub-populations performed equally well using the Predictive algorithm; however, only the sub-population which answered all of our recommendation comprehension questions correctly was able to further improve their performance using the Triage, Complementary, and Complementary Triage algorithms. Performing well with these algorithms required subjects to effectively parse when the recommendation was accurate or erroneous given their private information, thus we expect (and observe) that understanding how the algorithm generates recommendations is vital for its effective usage.
(a) Sub-population with complete comprehension
(b) Sub-population with comprehension errors
Figure 6: Average subject performance (through fractions of optimal hiring decisions) across treatments. $95\%$ bootstrap confidence intervals are clustered at the subject level. In addition, the triage performance (where compliance is perfect and not sending a recommendation uses the decision of Control) is given for each treatment by a gray dot, while the Response Effect from Section 3.2 is shown by the vertical red dashed lines. The optimal algorithm performance is shown by the gray horizontal line. The optimal performance achievable by a rational agent with the subject’s knowledge (role and algorithmic recommendation) is given for each treatment arm as a black dot.
Figure 7 reports how each sub-populations propensity to hire applications varied across the profiles they viewed. Generally subjects who answered at least one recommendation comprehension question wrong exhibited noisier decisions. This difference is particularly noticeable on Sales profiles for subjects in the Complementary and Complementary Triage treatments. Other than this excess noise, subjects who answered at least one recommendation comprehension question wrong responded to recommendations in a similar way to subjects who answered all recommendation comprehension questions correctly.
(a) Sub-population with complete comprehension
(b) Sub-population with comprehension errors
Figure 7: Frequency of experiment subjects taking a hire decision ($y$-axis) across the different combinations of treatments (color) and profiles ($x$-axis). Different markers are used to identify whether the subject received a hire recommendation ($\Uparrow$), did not receive a recommendation (--), or received a not-hire recommendation ($\Downarrow$). The optimal decision given oracle information is given by the location of the points on the lines $y=1$ (hire is optimal) and $y=0$ (not hire is optimal). The dashed lines give the frequency of an applicant being hired by a subject without access to recommendations (Control arm). The legend for these graphs is the same as the one in Figure 5 in the main text.